1. logx 8=3

2.log 10(2x-4) _ log 10(x+2) =1

Answer by stanbon(57222)

1. logx 8=3
---
x^3 = 8
x = 2
------------------
2.log 10(2x-4) - log 10(x+2) =1
---
log10[(2x-4)/(x+2)] = 1
---
[(2x-4)/(x+2)] = 10
----
2x-4 = 10x+20
8x = -24
x = -3
-----
But x cannot be -3 in the original equation
because you cannot have the log of a negative
number.
==============
No solution:
===================
Cheers,
stan H.